Reduce fraction to its lowest term
70/112

Found 2 solutions by fractalier, stanbon:
Answer by fractalier(2101)

70/112 = 35/56 = 5/8

Answer by stanbon(48545)

Reduce fraction to its lowest term
70/112
Divide numerator and denominator by "2" to get 35/56
Factor further to get [5*7]/[8*7]
Cancel the 7 factors to get 5/8
Cheers,
Stan H.